3) Consider preferences given by the utility function ?(?1, ?2) = ?1^2 + ?2. Derive the MRS and the individual demands for each of the two goods. How an increase in income would affect the demand of each good?
Answer #1

U = x12 + x2

Budget line: M = p1.x1 + p2.x2

MU1 = U/x1 = 2x1

MU2 = U/x2 = 1

MRS = MU1/MU2 = 2x1/1 = 2x1

Utility is maximized when MRS = p1/p2

2x1 = p1/p2

2p1.x1 = p2

p1.x1 = p2/2

x1 = (p2/2p1)

Substituting in budget line,

M = p1.x1 + p2.x2

M = (p2/2) + p2.x2

2M = p2 + 2p2.x2

2p2.x2 = 2M - p2

x2 = (2M - p2)/2p2

Demand function of x1 is independent of M, so increase in M does not affect demand for x1.

From demand function of x2, s M increases, (2M - p2) increases and so, demand for x2 increases.